8-pack control system working group Meeting Notes

8 July 2003


In attendance: R. Chestnut, R. Hall, K. Kim, J. Nelson, J. Rock

Note taker/typer: J. Nelson


Of immediate concern: What’s preventing the processing loop from working?


SWE is ready to test processing loop with real signals.  Sounds like there might not be time to test the loop until the next set of loads are installed.


Elias will send a crate profile to Ron/Judy/Kukhee of the crate he’s about to take a “spare” SIS out of.


Steve reported at the 10:00 8pack meeting that the LLRF calibration will be done in matlab.  Will need SWE support to add buttons to the EPICs panels to call these.


Kukhee reported that the vacuum readings are checked in the slow outer processing loop as a condition for turning back on after a trip.  High vacuum will not cause a trip.



- Janice will ask Elias if the SISes in TR09 have been modified as he put stickers on them (old item: Elias reported that this has been done: LLRF will make “MODIFIED” stickers for the SISes that have been modified very soon.)

- (still to do) Elias &/or Andrew will finish testing the SISes to make sure they work after the human error.

-Kukhee reported the database is done for the SISes in TR09.  He will give the PV name list ot Judy and Bob for displays and archiving respectively.

(done?) Kukhee will add an acquire button for these new SISes.

-this is done: the RMS calculations of AES SIS waveforms for the NLCTA are working.  Kukhee will add this feature for the 2 acoustic sensor SISes in TR09.

-(no change) We deferred the following: longer term issue: what to do if reading out an SIS when a fault occurs?



-About the DAC failure described below, Kukhee said getting status light status will take 1-2 days of work on the deep level code.  We decided to postpone that indefinitely.  He also reported that the values on our panels ARE the readback values, not the set values.

-Janice and Elias repoted that the DACs failed in an odd manner that was unsafe for the klystrons.  There was some discussion about modifying the driver.  Kukhee will look into the DAC manual and see if we can at least read back the module’s status (light on front is green or red?).  We put off asking for the DAC to check what was written to it for now.  Operations will have to decide.



-Still waiting on Cassel but we’re making a list...

-(no change) We’re currently waiting for Dick’s setup to settle down then we’ll do one more iteration of getting list, changing database, changing panels, programming the dcm, changing alarm handler and archive stuff as needed


DAC/ADC Issues

All of these issues are for sooner, Ron and Janice will try to get the details for implementation:


- The trickle down of the SLED variables.  This was specified at a previous meeting (numbers on SLED panel need to trickle down to I/Q and A/P PVs).  For processing, we might want this feature functional sooner rather than later. 

-The trickle down of the variables on the power distribution panel as well.  There are 4 PVs there called Klystron Sled-in phases.  These need to trickle down to the phase DACs as well.  This phase is to be added to the channel phase angle (IQ calib panel) and to the timestep phase before going out to the DAC.



-Janice will ask Andrew about:  Andrew will tell Kukhee when he takes out the DIO module to modify it for the deadman



-(I believe this is still true) Need to test the calibrated waveforms


Peak Power Meter (PPM)

-Janice reported that both PPMs are hooked up to labview, though it’s not running all the time during commissioning and load processing.

-Tonee has yet to write the labview code to retrieve the scope waveforms, but is working on it.

-Only the PPA PARMs were being archived.  Janice gave the new list to Judy/Bob.  Once the waveform retrieval is working, we’ll add those PVs to the archive list.

-Janice will teach llrf how to get at archived data and panel.



no changes to stuff below:


-(no change) On going: Need to check out readback with real signals, check archiving of variables (pwr, pmin, pmax, etc), and check the plots.



-(no change)Judy reported the crate profiles on the web are updated and look good.  Janice requested that channel assignment information also be available via the web, even it it’s just a link to the template file.  Judy and Ron will look into this.




Archiving of SISs

-(no change) SWE plans to implement the following two features requested by LLRF (Joe’s email):

have waveforms archived on (1) every trip and (2) when an acquire is done by a user

-(no change) Later SWE will work with Joe on his other two requests: archive every few minutes and when a change occurs in a selected set of values in EPICS (a pseudo-trip if you will)


Processing algorithm

(see above)

- Steve confirmed that LLRF folks will indeed contact SWE:  (still waiting on HW folks) Processing algorithm ready to be checked.  Andy will call Kukhee and Ron when there’s a good time to check it.

- Kukhee will make sure the processing loop resets the IQAs too.

-The trigger stopping mechanism on a fault was discussed.  Joe proposed a DIO bit goes to NIM logic etc etc.  It’s believed that Kukhee’s method of changing the trigger register status is too slow for the number of SISes and their location in different crates.